OpenStudy (anonymous):

I takes 3.5 hours to drive to davenport to hogsdale and then 2 hours and 10 minutes to drive from hogsdale to paqua. On the route that she took it is 720 miles from davenport to paqua. What is her average driving speed for the trip

OpenStudy (anonymous):

3.5 hrs + 2 hrs + 10 minutes = 5 hours and 40 minutes = 5 hours and 2/3 hours Average speed = s / t =720 miles / (5 hours + 2/3 hours) =720 miles / (17/3 hours) = 127.06 miles/hours (approximation)